OpenStudy (bekkah323):

2x+7/x-6 plus 3x/6-x Equation below

OpenStudy (bekkah323):

\[\frac{ 2x+7 }{ x-6 } + \frac{ 3x }{ 6-x }\]

OpenStudy (marissalovescats):

Well.. Your denominators are the same just switched around. So all you have to do is add the numerators.

OpenStudy (marissalovescats):

If I'm not mistaken that is.

OpenStudy (bekkah323):

don't i have a -6 and a regular 6 though?

OpenStudy (jdoe0001):

6-x => -(x-6)

jimthompson5910 (jim_thompson5910):

marissalovescats, you're close, but you would factor 6-x to get -1(x-6) so \[\large \frac{2x+7}{x-6} + \frac{3x}{6-x}\] \[\large \frac{2x+7}{x-6} + \frac{3x}{-1(x-6)}\] \[\large \frac{2x+7}{x-6} - \frac{3x}{x-6}\] now you can combine the fractions

OpenStudy (marissalovescats):

Ah yes. Truth ^

OpenStudy (bekkah323):

I'm a little confused on what you did?

jimthompson5910 (jim_thompson5910):

do you see how 6-x = -1(x-6) ?

OpenStudy (bekkah323):

if you distribute that doesn't it make -x+6?

jimthompson5910 (jim_thompson5910):

then you rearrange terms to get 6 - x

OpenStudy (bekkah323):

oh, ok

jimthompson5910 (jim_thompson5910):

6 - x = 6 + (-x) 6 - x = -x + 6 6 - x = -1*x + (-1)*(-6) 6 - x = -1*(x - 6)

jimthompson5910 (jim_thompson5910):

do you see how we went from \[\large \frac{2x+7}{x-6} + \frac{3x}{-1(x-6)}\] to \[\large \frac{2x+7}{x-6} - \frac{3x}{x-6}\]

OpenStudy (bekkah323):

yes

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (bekkah323):

thanks

jimthompson5910 (jim_thompson5910):

sure thing